OpenStudy (anonymous):

cosx=5/13 and sin y=3/5 what is the exact value of sin x?

OpenStudy (anonymous):

sin x = 12/13

OpenStudy (anonymous):

one more method to do it is to keep in mind the identity \[\cos ^{2}x+\sin ^{2}x =1 \]

OpenStudy (anonymous):

so (5/13)^2 + sin^2x = 1 sin^2x = 1-25/169 = 144/169 so sin x = sqrt (144/169) = 12/13

OpenStudy (anonymous):

so what was sin y for?

OpenStudy (anonymous):

for nothing? lol